for the reaction 2NH3 equilibrium arrows 3H2 + N2, the value of K at 500 celsius is 17. calculate the value of Kp for this reaction at 500 C
Answer #1

N2(g) + 3H2(g) ...............> 2NH3(g)

Dn = 2- (1+3) = -2

T = 500 +273 = 773

Kp = Kc*(RT)^Dn

R = universal gaseous constant = 0.0821 lit.atm/mol*k

Kp = 17*(0.0821*773)^(-2)

Kp = 4.2*10^-3 atm^-2